The difference between 2 vectors has a magnitude equal to the magnitude of either vector. The angle between two vectors is

Grade:12

1 Answers

Sanju
6 years ago
If we assume vectors as X, then the magnitude of the resultant vector is also X. This is mentioned in the question itself.
To find the resultant vector, we have an equation: R2 = A2+B2+2ABcosθ
Applying the above formula gives: X2 = X2+X2+2XXcosθ
X2 = 2X2+2X2cosθ
-X= 2X2cosθ
-1/2 = cosθ
Therefore, θ = cos-1(-1/2) = 120
ANSWER: ANGLE BETWEEN 2 VECTORS IS 120 DEGREE